Evaluate 𝐹 βˆ™ 𝑛𝑑𝑆, where 𝐹 = 4π‘₯𝑧𝑖 βˆ’ 𝑦 2 𝑗 + π‘¦π‘§π‘˜ and S is the surface of the bounded by π‘₯ = 0, π‘₯ = 1, 𝑦 = 0, 𝑦 = 1, 𝑧 = 0, 𝑧 = 1.


1
Expert's answer
2021-05-26T14:09:05-0400

The Divergence Theorem


"\iint\vec F\cdot d\vec S=\iiint \text{div}\vec F dV""S\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ V \ \ \ \ \ \ \ \ \ \ \ \ \"

The divergence of "\vec F" is


"\text{div}\vec F=\nabla\cdot \vec F"

"=(i\dfrac{\partial}{\partial x}+j\dfrac{\partial}{\partial y}+k\dfrac{\partial}{\partial z})(4xzi− y^2j + yzk)"

"=\dfrac{\partial}{\partial x}(4xz)+\dfrac{\partial}{\partial y}(-y^2)+\dfrac{\partial}{\partial z}(yz)=4z-2y+y"

"=4z-y"

Then


"\iint\vec F\cdot d\vec S=\iiint \text{div}\vec F dV""S\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ V \ \ \ \ \ \ \ \ \ \ \ \ \"

"=\displaystyle\int_{x=0}^{1}\displaystyle\int_{y=0}^{1}\displaystyle\int_{z=0}^{1}(4z-y)dzdydx"

"=\displaystyle\int_{x=0}^{1}\displaystyle\int_{y=0}^{1}[2z^2-yz]\begin{matrix}
 1\\
 0
\end{matrix}dydx"

"=\displaystyle\int_{x=0}^{1}\displaystyle\int_{y=0}^{1}(2-y)dydx"

"=\displaystyle\int_{x=0}^{1}[2y-\dfrac{y^2}{2}]\begin{matrix}
 1 \\
 0
\end{matrix}dx=\displaystyle\int_{x=0}^{1}\dfrac{3}{2}dx"

"=\dfrac{3}{2}[x]\begin{matrix}
 1 \\
 0
\end{matrix}=\dfrac{3}{2}"



"\iint\vec F\cdot d\vec S=\dfrac{3}{2}""S\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \"